What is the total energy of an electron when its momentum is 10 times its rest mass multiplied by the light velocity?
1
Expert's answer
2019-05-09T11:31:42-0400

For this problem we use formula

"E=c\sqrt{m_0^2c^2+p^2}"

initial condition

"p=10m_0c"

then

"E=m_0c^2\sqrt{11}"

rest energy for electron

"E_0=m_0c^2=0.511 (MeV)"

then

"E=E_0\sqrt{11}=0.511*\sqrt{11}=1.695 (MeV)"

In joules it

"E=1.695 \cdot 10^6 \cdot 1.6 \cdot 10^{-19}=2.711\cdot 10^{-13} (J)"